A full journal bearing of 50mm diameter and 100mm long has a bearing pressure of 1.4N/mm2
. The
speed of the journal bearing is 900r.p.m and the ratio of journal diameter to the diametral clearance is
1000. The bearing is lubricated with oil whose absolute viscosity at the operating temperature of 75oC. Find: (a) The amount of artificial cooling required, and (b) The mass of the lubricating oil required, if the
difference between the outlet and inlet temperature of the oil is 10oC. Take specific heat of the oil as
1850J/kg/ oC.
1
Expert's answer
2020-05-04T01:53:40-0400

For finding the artificial cooling we need to find the coeffciient of friction


"\mu=(\frac{33}{10^8})\times(\frac{ZN}{P})(\frac{d}{c})+k=(\frac{33}{10^8})\times(\frac{0.011\times900}{1.4})(1000)+0.002=0.004"

Artificial cooling= Heat generated- Heat dissipated


Artificial cooling ="(\mu W V)-(Cld\Delta T)"


Artificial cooling="(0.004\times (1.4\times 50\times100)\times2.36)-(280\times0.05\times0.1\times 10)"


Artificial cooling= 52.08 J/s

(b) mass of lubricating oil required= mc "\Delta T"

52.08"=m\times 1850 \times 10"


m="2.81\times 10^{-3} kg /s"